OpenStudy (anonymous):

a proton of mass 1.67x10^-27 kg is travelling with an ititial speed of 1.57km/s when is collides head on with a stationary alpha particle of mass 6.64x10^-27kg. the proton rebounds with a speed of 0.893km/s. what is the speed of the alpha particle?

OpenStudy (experimentx):

use conservation of linear momentum mas of alpha particle is around 4 times the mass of proton.

OpenStudy (experimentx):

m1u1 + m2 0 = m1v1 + m2v2 v2 = m1(u1 - v1)/m2

OpenStudy (anonymous):

what is u1?

OpenStudy (experimentx):

velocity of proton

OpenStudy (anonymous):

i get 0.170 and the answer is 0.619?

OpenStudy (mos1635):

u1 = initial velocity=1.57km/s v1=final velocity = -0.893km/s !!!!!!!

OpenStudy (anonymous):

else use conservation of kinetic energy,if its an elatic collision....

OpenStudy (anonymous):

negative sign! right.... :S got it thanks!